OpenStudy (anonymous):

y − 4 =3/4 (x + 8) , identify the equation of the same line in standard form.

OpenStudy (amorfide):

well first of all, multiply both sides by 4 to get rid of the fraction 4(y-4)=3(x+8) multiply out the brackets on both sides 4y-16=3x+24 add 17 to both sides 4y=3x+40 divide by 4 on both sides to get y on its own